OpenStudy (anonymous):

why does 9/x^2 -1 simplify to 9-x^2/x^2 . is it because we need to reduce the expression to one term?

OpenStudy (datanewb):

\[\frac{9}{x^2} - 1 = \frac{9-x^2}{x^2} \] Notice that 1 can be rewritten as \[\frac{x^2}{x^2}\] You can try any value of for x, and it will always reduce to 1/1 = 1. So then we have: \[\frac{9}{x^2} - \frac{x^2}{x^2} \] and since those have common denominators, we can just subtract the numerators to get \[\frac{9-x^2}{x^2} \]

OpenStudy (anonymous):

So are we ending up with an expression that is supposed to be simpler than the original, and why?

OpenStudy (datanewb):

Oh, sorry, I see what you are asking. I think it depends on the situation as to which side is simpler. I guess the right side might be simpler for computation, because you can simply input values for x when calculating it's value without then having to do the extra step of finding a common denominator, but for the most part I find both sides equally simple/complex.

OpenStudy (anonymous):

Ok, thanks, I thought it was cast in stone that it had to be simplified. Your answer really helps!